# Tensor Product Attention Is All You Need

### v3

@@ A.6 Overall Complexity for Single-Head
    QK $\mathbf{b}$-Dot Product Stage: $\mathcal{O}(T^{2}\,R_{q}\,R_{k}\,d_{h})$.
@@
    QK Scalar-Multiply Stage: $\mathcal{O}(T^{2}\,R_{q}\,R_{k})$.
@@
    Computing $\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})$ for all tokens: $\mathcal{O}(T\,R_{v}\,d_{h})$.
  4. (iv)
    Weighted Summation by $\alpha_{t,\tau}\,a_{v}^{(u)}(\mathbf{x}_{\tau})$: $\mathcal{O}(T^{2}\,R_{v}\,d_{h})$.
Hence, for a single head, the total cost is:
$\mathcal{O}\Bigl{(}T^{2}\,R_{q}\,R_{k}\,d_{h}~{}+~{}T^{2}\,R_{q}\,R_{k}~{}+~{}T\,R_{v}\,d_{h}~{}+~{}T^{2}\,R_{v}\,d_{h}\Bigr{)}.$
In many cases (especially for large $T$), the $\mathcal{O}(T^{2})$ terms dominate, so one often focuses on
$\mathcal{O}\Bigl{(}T^{2}\,R_{q}\,R_{k}\,d_{h}~{}+~{}T^{2}\,R_{q}\,R_{k}~{}+~{}T^{2}\,R_{v}\,d_{h}\Bigr{)}.$
A.7 Multi-Head and Batch Extensions (Reuse of $\mathbf{b}$-Dot Products)
When extending to $h$ heads and batch size $B$, all sequence‐length‐dependent terms are multiplied by $\sim B\,h$. However, crucial $\mathbf{b}$-dot products can be shared across heads:
QK $\mathbf{b}$-Dot Products. Since each head has distinct scalar factors $a_{q,i},\,a_{k,i}$ but the same $\mathbf{b}_{q}^{(r)},\,\mathbf{b}_{k}^{(s)}$ across heads, each pairwise dot product
$\langle\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),\,\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})\rangle$
is computed just once per batch. That cost remains
$\mathcal{O}\bigl{(}B\,T^{2}\,R_{q}\,R_{k}\,d_{h}\bigr{)},$
not multiplied by $h$. After caching these dot products, each of the $h$ heads pays $\mathcal{O}\bigl{(}B\,T^{2}\,h\,R_{q}\,R_{k}\bigr{)}$ total for the head‐specific scalar multiplications (the “$\alpha_{t,\tau}$”–like factors).
V $\mathbf{b}$-Evaluations. Likewise, the $\mathbf{b}_{v}^{(u)}$ factors are shared across heads (i.e. one set of $\mathbf{b}_{v}$-vectors for all heads). Hence, computing all $\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})$ for $\tau=1\ldots T$ (across the batch) is a one‐time cost:
$\mathcal{O}\bigl{(}B\,T\,R_{v}\,d_{h}\bigr{)}.$
Then each head $i$ has its own scalar factors $a_{v,i}^{(u)}(\mathbf{x}_{\tau})$, so the final accumulation $\sum_{\tau=1}^{T}\alpha_{t,\tau}\,a_{v,i}^{(u)}(\mathbf{x}_{\tau})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})$ costs $\mathcal{O}\bigl{(}B\,T^{2}\,h\,R_{v}\,d_{h}\bigr{)}$ in total (for all $t,u$).
Putting it all together, the total flops for multi‐head attention with batch size $B$ are:
@@ A.7 Multi-Head and Batch Extensions (Reuse of $\mathbf{b}$-Dot Products)
$\underbrace{\mathcal{O}\bigl{(}B\,T^{2}\,R_{q}\,R_{k}\,d_{h}\bigr{)}}_{\begin{subarray}{c}\text{QK $\mathbf{b}$-dot products}\\ \text{(shared across heads)}\end{subarray}}~{}+~{}\underbrace{\mathcal{O}\bigl{(}B\,T^{2}\,h\,R_{q}\,R_{k}\bigr{)}}_{\text{per-head QK scalar mult.}}~{}+~{}\underbrace{\mathcal{O}\bigl{(}B\,T\,R_{v}\,d_{h}\bigr{)}}_{\begin{subarray}{c}\text{Compute $\mathbf{b}_{v}$ for all tokens}\\ \text{(shared across heads)}\end{subarray}}~{}+~{}\underbrace{\mathcal{O}\bigl{(}B\,T^{2}\,h\,R_{v}\,d_{h}\bigr{)}}_{\begin{subarray}{c}\text{final accumulations}\\ \text{(per head)}\end{subarray}}.$
Discussion. By contrast, standard multi-head attention typically requires $\mathcal{O}\bigl{(}B\,T^{2}\,h\,d_{h}\bigr{)}$ flops for the $\mathbf{Q}\mathbf{K}^{\top}$ dot product (plus a similar $\mathcal{O}\bigl{(}B\,T^{2}\,h\,d_{h}\bigr{)}$ for multiplying by $\mathbf{V}$). The factorization can yield savings provided $R_{q}R_{k}\ll h$ (for QK) and $R_{v}\ll h$ (for V), though actual speedups depend on how well these multi‐stage kernels are implemented and on hardware efficiency. By retaining $\mathbf{Q},\mathbf{K},$ and $\mathbf{V}$ in factorized form, one can forgo the usual steps:
@@
$\mathbf{x}_{t}~{}\mapsto~{}\mathbf{Q}_{t},~{}\mathbf{K}_{\tau}~{}\mapsto~{}(\mathbf{Q}\,\mathbf{K}^{\top})~{}\mapsto~{}\operatorname{softmax}(\mathbf{Q}\,\mathbf{K}^{\top})\,\mathbf{V}~{}\mapsto~{}\text{final output}.$
Instead, the large $\mathbf{Q},\mathbf{K},\mathbf{V}$ tensors (of size $T\times d_{h}$) are never materialized. The cost is replaced by rank-based $\mathbf{b}$-dot‐product computations plus per‐head scalar multiplications. The main challenge is to keep the factor ranks $(R_{q},R_{k},R_{v})$ sufficiently small relative to $d_{h}$ and to implement the necessary multi‐stage kernels efficiently. When $R_{q},R_{k},R_{v}\ll h$, fully factorized QKV attention can yield substantial gains in both computation and memory footprint.
A.8 Decoding Speed during Inference Time of MHA, MQA, GQA, MLA, and TPA
Suppose we are in an autoregressive setting, decoding the current token $\mathbf{x}_{T}$ given cached keys and values (KV) from all previous tokens $\mathbf{x}_{1},\dots,\mathbf{x}_{T-1}$. For each attention head $i\in\{1,\dots,h\}$, we store $\mathbf{K}_{i}\in\mathbb{R}^{T\times d_{h}},\mathbf{V}_{i}\in\mathbb{R}^{T\times d_{h}}$. Below, we compare the flops needed by MHA, MQA, GQA, MLA, and TPA to compute the next-token logits during inference.
MHA, MQA, and GQA. Despite sharing or grouping keys/values in MQA and GQA, the decoding cost for MHA, MQA, and GQA remains of the same order. Specifically, for each head $i$, we compute:
$\mathbf{Q}_{i}(\mathbf{x}_{T})~{}\in~{}\mathbb{R}^{d_{h}},\quad\mathbf{K}_{i}~{}\in~{}\mathbb{R}^{T\times d_{h}},\quad\mathbf{Q}_{i}(\mathbf{x}_{T})\,\mathbf{K}_{i}^{\top}~{}\in~{}\mathbb{R}^{1\times T},$
$\quad\text{and}\quad\operatorname{Softmax}\bigl{(}\mathbf{Q}_{i}(\mathbf{x}_{T})\,\mathbf{K}_{i}^{\top}\bigr{)}\,\mathbf{V}_{i}~{}\in~{}\mathbb{R}^{d_{h}}.$
Hence, the flops scale linearly in $h$, $d_{h}$, and $T$. For example, forming $\mathbf{Q}_{i}(\mathbf{x}_{T})\mathbf{K}_{i}^{\top}$ for each head $i$ costs roughly $\mathcal{O}(h\,d_{h}\,T)$.
MLA. During inference, MLA can be seen as MQA but uses a larger head dimension to accommodate both RoPE and compressed representations (e.g., $d_{h}^{\prime}=d_{\text{rope}}+d_{c}$). In typical configurations, $d_{\text{rope}}+d_{c}$ can be significantly larger (e.g., $d_{h}^{\prime}=576$ rather than $d_{h}=64$ or $128$), thus inflating the dot-product cost by roughly 4.5×4.5$\times$4.5 × to 9×9$\times$9 × compared to MHA/MQA/GQA.
TPA. Recall that TPA factorizes $\mathbf{Q}$ and $\bm{K}$ into rank-$(R_{q},R_{k})$ terms (see Section A), potentially avoiding large $\mathbf{Q},\bm{K}$ materializations. At inference, TPA’s dot-product cost can be broken into two parts:
$\underbrace{R_{q}\,R_{k}\,d_{h}\,T}_{\text{QK $\mathbf{b}$-dot products (shared across all heads)}}~{}+~{}2\,\underbrace{R_{q}\,R_{k}\,h\,T}_{\text{per-head scalar multiplications}},$
where $T$ is the current sequence length. For concrete values $d_{h}=128$, $h=64$, $R_{q}=8$, and $R_{k}=2$ (or $R_{q}=16$, $R_{k}=1$), we obtain:

### v4

@@ C.6 Overall Complexity for Single-Head (Specialized)
    QK $\mathbf{B}$-Dot Product Stage ($P$): $\mathcal{O}(T^{2}\,R_{Q}\,R_{K}\,d_{h})$ (shared calculation if multi-head).
@@
    QK $\mathbf{A}$-Factor Combination and Scaling Stage ($\mathcal{L}^{\prime}$): $\mathcal{O}(T^{2}\,R_{Q}\,R_{K})$. (Scaling by $s_{Q}s_{K}$ is part of this per-element operation on scores).
@@
    Value Aggregation Stage (including $s_{V}$ scaling): $\mathcal{O}(T^{2}\,R_{V}\,d_{v})$.
Total for single head, assuming $P$ is computed once: $\mathcal{O}\Bigl{(}T^{2}\,R_{Q}\,R_{K}\,d_{h}+T^{2}\,R_{Q}\,R_{K}+T^{2}\,R_{V}\,d_{v}\Bigr{)}$.
C.7 Multi-Head and Batch Extensions (Reuse of $\mathbf{B}$-Dot Products)
@@
QK $\mathbf{B}$-Dot Products ($P$): Shared across heads. Cost: $\mathcal{O}\bigl{(}B\,T^{2}\,R_{Q}\,R_{K}\,d_{h}\bigr{)}$.
Per-Head QK $\mathbf{A}$-Factor Combination and Scaling ($\mathcal{L}^{\prime}$): Each head combines its $\mathbf{A}_{Q},\mathbf{A}_{K}$ with $P$ and applies $s_{Q}s_{K}$. Cost: $\mathcal{O}\bigl{(}B\,H\,T^{2}\,R_{Q}\,R_{K}\bigr{)}$.
Value Aggregation (including $s_{V}$): Assuming $\mathbf{B}_{V}$ factors are shared and $\mathbf{A}_{V}$ are per-head. Cost: $\mathcal{O}\bigl{(}B\,H\,T^{2}\,R_{V}\,d_{h}\bigr{)}$. Assuming $d_{v}=d_{h}$. Total FLOPs for specialized multi-head TPA (full sequence processing) with batch size $B$:
$\mathcal{F}_{\text{TPA-specialized}}=\underbrace{\mathcal{O}\bigl{(}B\,T^{2}\,R_{Q}\,R_{K}\,d_{h}\bigr{)}}_{\text{QK $\mathbf{B}$-dot products }P}+\underbrace{\mathcal{O}\bigl{(}B\,H\,T^{2}\,R_{Q}\,R_{K}\bigr{)}}_{\text{Per-head QK }\mathbf{A}\text{-comb. + }s_{Q}s_{K}\text{ scale}}+\underbrace{\mathcal{O}\bigl{(}B\,H\,T^{2}\,R_{V}\,d_{h}\bigr{)}}_{\text{Value Aggregation+ }s_{V}\text{ scale}}.$
Discussion. By contrast, standard multi-head attention typically requires $\mathcal{F}_{\text{MHA}}=\mathcal{O}\bigl{(}2B\,T^{2}\,H\,d_{h}\bigr{)}$ FLOPs (for $\mathbf{Q}\mathbf{K}^{\top}$ and attention-weighted $\mathbf{V}$). The specialized TPA can be more efficient if $R_{Q}R_{K}d_{h}+HR_{Q}R_{K}+HR_{V}d_{h}<2Hd_{h}$. Dividing by $Hd_{h}$, this condition becomes $(R_{Q}R_{K}/H)+(R_{Q}R_{K}/d_{h})+R_{V}<2$. For example, if $R_{Q}=R_{K}=R_{V}=1$, the condition is $21/H+1/d_{h}+1<$ 2, or $11/H+1/d_{h}<$ 1, which is generally true for typical $H,d_{h}\geq 2$. Thus, with small ranks, TPA can offer computational savings. Actual speedups also depend critically on memory access patterns and kernel implementations, as demonstrated by FlashAttention-style approaches. By retaining $\mathbf{Q},\mathbf{K},$ and $\mathbf{V}$ in factorized form, one can bypass the explicit materialization of these large tensors:
@@
$\mathbf{x}_{t}\rightarrow\mathbf{Q}_{t},~{}\mathbf{K}_{\tau},\mathbf{V}_{\tau}\quad(\text{materialized})\rightarrow(\mathbf{Q}\,\mathbf{K}^{\top})\rightarrow\operatorname{softmax}(\mathbf{Q}\,\mathbf{K}^{\top})\,\mathbf{V}\rightarrow\text{final output}.$
Instead, the large $\mathbf{Q},\mathbf{K},\mathbf{V}$ tensors (each of size $B\times T\times H\times d_{h}$) are not explicitly formed. The computation is restructured into rank-based $\mathbf{B}$-dot-product computations, per-head $\mathbf{A}$-factor combinations, and appropriate scaling. The challenge lies in choosing ranks $(R_{Q},R_{K},R_{V})$ small enough to ensure computational benefits while maintaining model quality, and in implementing the multi-stage kernels efficiently. When ranks are sufficiently small, this specialized approach can lead to gains in both computation and memory footprint.
@@
Figure 6: Data flow diagram for specialized TPA computation, avoiding materialization of full $\mathbf{Q},\mathbf{K},\mathbf{V}$ tensors. Phase 1 (Top): Attention Score Computation. Factorized $\mathbf{B}_{Q},\mathbf{B}_{K}$ produce shared dot-products $P$ (summing over feature dimension $d$). $P$ is combined with factorized $\mathbf{A}_{Q},\mathbf{A}_{K}$ and scaled by $s_{Q}s_{K}$ (summing over ranks $r_{q},r_{k}$) to form scaled logits $\mathcal{L}^{\prime}$. These are then scaled by $1/\sqrt{d_{h}}$ and passed through Softmax to yield attention probabilities $\bm{\alpha}$. Phase 2 (Bottom): Value Aggregation. $\bm{\alpha}$ weights the factorized $\mathbf{A}_{V}$ (conceptually forming $W_{\mathbf{A}_{V}}$ via an einsum-like combination), which then aggregates with $\mathbf{B}_{V}$ and is scaled by $s_{V}$ (summing over key tokens $t_{k}$ and value ranks $r_{v}$) to produce the final output $\mathbf{O}$. (Diagram simplifies batch dimensions $B$; $T$ denotes sequence length; $H$ is heads; $R_{X}$ are ranks $R_{Q},R_{K},R_{V}$ corresponding to indices $r_{q},r_{k},r_{v}$; $d_{h}$ is the feature dimension $d$ for $\mathbf{B}$ factors. Operations $\sum_{X}$ denote contractions over index $X$, and $\odot$ denotes an einsum-like combination. Scaling factors $s_{Q},s_{K},s_{V}$ are incorporated as described.).
C.8 Inference Time Decoding of Different Attention Mechanisms
In autoregressive decoding, we generate the output for the current token $\mathbf{x}_{M}$ (where $M$ is the current sequence length, previously denoted $T$ in this section for a generic single token, but $M$ is used in Alg. 1) given cached keys and values from $M-1$ previous tokens. We analyze the FLOPs for computing the attention output for this single query token. For all mechanisms, we consider the FLOPs for computing scaled logits and aggregating values, excluding the initial projection of the current token’s hidden state $\mathbf{x}_{M}$ into Q, K, V representations.
MHA, MQA, and GQA. For Multi-Head Attention (MHA), with $H$ query heads and $H$ distinct Key/Value heads:
  • Logits: Each of the $H$ query vectors (dim $d_{h}$) interacts with its corresponding K-cache (size $M\times d_{h}$). Cost: $H\cdot(M\cdot d_{h})=MHd_{h}$.
  • Value Aggregation: Each of the $H$ attention patterns interacts with its V-cache (size $M\times d_{h}$). Cost: $H\cdot(M\cdot d_{h})=MHd_{h}$.
  • Total MHA: $2MHd_{h}$.
Multi-Query Attention (MQA) uses $H$ query heads but shares a single Key/Value head ($H_{kv}=1$).
  • Logits: Each of the $H$ query vectors (dim $d_{h}$) interacts with the single shared K-cache (size $M\times d_{h}$). Cost: $H\cdot(M\cdot d_{h})=MHd_{h}$.
  • Value Aggregation: Each of the $H$ attention patterns interacts with the single shared V-cache (size $M\times d_{h}$). Cost: $H\cdot(M\cdot d_{h})=MHd_{h}$.
  • Total MQA: $2MHd_{h}$.
Grouped-Query Attention (GQA) uses $H$ query heads and $N_{g}$ Key/Value head groups ($H_{kv}=N_{g}$). The $H$ query heads are partitioned into $N_{g}$ groups, each containing $H/N_{g}$ query heads that share one K/V head pair within that group.
  • Logits: For each of the $N_{g}$ groups, $H/N_{g}$ query vectors interact with the group’s K-cache. Cost per group: $(H/N_{g})\cdot(M\cdot d_{h})$. Total for $N_{g}$ groups: $N_{g}\cdot(H/N_{g})\cdot Md_{h}=MHd_{h}$.
  • Value Aggregation: Similarly, $MHd_{h}$.
  • Total GQA: $2MHd_{h}$.
MQA and GQA significantly reduce the KV cache size and memory bandwidth requirements compared to MHA. While the arithmetic FLOP count for the core attention computation (dot products and weighted sums) is $2MHd_{h}$ for all three if they have the same number of query heads $H$ and head dimension $d_{h}$, the practical speedups for MQA/GQA often arise from better memory access patterns due to smaller K/V caches.
MLA. Multi-Head Latent Attention (MLA), as described in Appendix H.3, uses $H$ heads. For each head, the key $\mathbf{K}_{i}$ has dimension $d_{k}^{\prime}=d_{c}+d_{h}^{R}$ (compressed part + RoPE part), and the value $\mathbf{V}_{i}^{C}$ has dimension $d_{v}^{\prime}=d_{c}$ (only compressed part).
  • Logits: $H$ query heads (dim $d_{k}^{\prime}$) interact with K-caches (dim $d_{k}^{\prime}$). Cost: $MHd_{k}^{\prime}$.
  • Value Aggregation: $H$ attention patterns interact with V-caches (dim $d_{v}^{\prime}$). Cost: $MHd_{v}^{\prime}$.
  • Total MLA: $MH(d_{k}^{\prime}+d_{v}^{\prime})=MH(d_{c}+d_{h}^{R}+d_{c})=MH(2d_{c}+d_{h}^{R})$.
TPA. We use the FlashTPA Decoding algorithm (Algorithm 1) for FLOPs analysis, with $N=1$ query token, $M$ cached items, $D$ as feature dimension for $\mathbf{B}_{Q}/\bm{b}^{K}$ (typically $d_{h}$), and $E$ for $\bm{b}^{V}$ (typically $d_{h}$). For ranks $(R_{Q},R_{K},R_{V})$:
  • Logits computation (Steps 1-3 of Alg. 1): $\mathcal{O}(M(R_{Q}R_{K}D+HR_{Q}R_{K}+HR_{K}))$.
  • Value aggregation (Steps 5-6 of Alg. 1): $\mathcal{O}(M(HR_{V}+HR_{V}E))$.
  • Total for TPA decoding: $\mathcal{O}(M[R_{K}(R_{Q}D+HR_{Q}+H)+R_{V}H(1+E)])$.
Example Comparison.
Let’s use the ranks from FlashTPA experiments (Section 5.2): $(R_{Q},R_{K},R_{V})=(16,1,1)$, and typical dimensions $D=E=d_{h}=64$, $H=32$ (e.g., for a 2048 $d_{\text{model}}$).
MHA Logits: $\quad M\cdot H\cdot d_{h}=M\cdot 32\cdot 64=2048M$
MHA Value Agg.: $\quad M\cdot H\cdot d_{h}=M\cdot 32\cdot 64=2048M$
MHA Total: $\quad 4096M$
TPA $\text{TPA Logits}(R_{K}=1,R_{Q}=16):$ $\quad M(R_{Q}d_{h}+HR_{Q}+H)=M(16\cdot 64+32\cdot 16+32)$
$=M(1024+512+32)=1568M$